OpenStudy (anonymous):

Consider the line y=-9x+5 . Find the equation of the line that is perpendicular to this line and passes through the point −8, 4 . Find the equation of the line that is parallel to this line and passes through the point −8, 4 .

OpenStudy (campbell_st):

ok... so perpendicular lines have a slope that is the negative reciprocal of the given slope parallel lines have the same slope so you need to start by stating the slope of the given line what do you think..?

OpenStudy (anonymous):

m = -9 and b= 5

OpenStudy (campbell_st):

ok... so we only need m so the negative recpirocal, for the perpendicular line is 1/9 so the slope of the perpendicular line is m = 1/9 and you have a point (-8, 4) so you can find the equation. the parallel line has the same slope as the original line so m = -9 and you have point (-8, 4) I tend to short cut this perpendicular is y = 1/9 x + b substitute x = -8 and y = 4 to find b parallel same thing y = -9x + b substitute x = -8 and y = 4 to find b hope it helps

OpenStudy (anonymous):

yes thank you so much!